OpenStudy (anonymous):

Find the solutions to the system. y=x^2+5x+6 y=4x+12 Please help me..

OpenStudy (anonymous):

.........Anybody ??

OpenStudy (thefurball):

set the two equations equal to each other and factor: 4x + 12 = x^2 + 5x + 6 0 = x^2 + x - 6 0 = (x+3)(x-2) x=-3,x=2 plug in the two points y = -12 + 12 = 0 y = 8 + 12 = 20 therefore the two points are (-3,0) and (2,20)